# ] 1NLPR, MAIS, CASIA  2UCAS  3ByteDance

### v1

@@ 5 Experiments
Figure 11: Performance decoupling with AI2D [17].
Figure 12: Performance decoupling with MathVista [37].
Figure 13: Performance decoupling with MMStar [7].
5.1 Main Properties on TreeBench
Results on TreeBench. We evaluate four state-of-the-art private models, including GPT-4o-1120 [41] and o3-0416 [43] from OpenAI, and Gemini-2.5-Flash-0520 [9] and Gemini-2.5-Pro-0605 [10] from Google DeepMind. Additionally, 8 representative open-source general models are incorporated, including LLaVA-OneVision series [23], Qwen2.5-VL series [2], and InternVL3 series [77]. Furthermore, two very recent visual grounded reasoning models are also included, i.e., DeepEyes [76] and Pixel-Reasoner [50], as both of them follow a “grounding then answering” pipeline, with the capability of “thinking with images”. Table 2 presents per per-category performance of different models. Overall, OpenAI’s o3-0416 [43], the state-of-the-art visual grounded reasoning model, demonstrates the strongest perception abilities, as expected. Larger models usually contribute to better performance, while Qwen2.5-VL-32B [2] and its 72B counterpart as an exception. Notably, our TreeVGR-7B even achieves comparable performance with InternVL3-78B [77], demonstrating the effectiveness of the visual grounded reasoning pipeline.
Correlation between Localization and Performance. Importantly, for visual grounded reasoning models, our traceable evaluation demonstrates a positive correlation between localization precision and the overall performance, as illustrated in Table 2. This positive correlation between precise localization (mIoU) and overall performance is evident in the progressive improvement from DeepEyes-7B [76] to Pixel-Reasoner-7B [50] to our final TreeVGR-7B. As mIoU increases, the overall scores rise correspondingly, with TreeVGR-7B achieving the highest mIoU and strongest overall performance at the same time.
@@ 5.1 Main Properties on TreeBench
V* Bench [66] HR-Bench-4K [63] HR-Bench-8K [63]
Overall Attr. Spatial Overall Single Cross Overall Single Cross
Private Models
GPT-4o-1120 [41] 66.0
o3-0416 [43] 95.7
Open-source General Models
LLaVA-OneVision-7B [23] 70.7 73.0 60.5 64.3 74.8 53.8 59.8 65.3 54.3
LLaVA-OneVision-72B [23] 73.8 80.9 63.2 66.3 76.5 56.0 60.9 68.8 53.0
InternVL3-8B [77] 72.3 73.0 71.1 70.8 79.3 62.3 62.0 64.3 59.8
InternVL3-38B [77] 77.5 77.4 77.6 76.3 83.5 69.0 67.0 71.3 62.8
InternVL3-78B [77] 76.4 75.7 77.6 75.5 84.5 66.5 67.3 71.8 62.8
Qwen2.5-VL-7B [2] 74.3 77.4 69.7 72.1 88.8 55.5 68.8 83.5 54.0
Qwen2.5-VL-32B [2] 85.9 83.5 89.5 74.8 89.3 60.3 71.6 86.5 56.8
Qwen2.5-VL-72B [2] 84.8 90.8 80.9 79.4 88.8 70.0 76.3 84.3 68.3
Open-source Visual Grounded Reasoning Models
Pixel-Reasoner-7B [50] 80.6 83.5 76.3 72.9 86.0 60.3 66.9 80.0 54.3
DeepEyes-7B [76] 90.0 92.1 86.8 75.1 91.3 59.0 72.6 86.8 58.5
TreeVGR-7B 91.1 94.0 87.0 77.1 90.3 64.0 73.1 86.5 59.8
Table 3: Comparison with state-of-the-art alternatives on V* Bench [66] and HRBench [63]. All results are self-collected. Best performances of visual grounded reasoning models are highlighted in bold.
Perception Reasoning
Overall OCR RS DT MO AD OCR DT MO AD
General Models
Qwen2.5-VL-7B [2] 42.3 87.6 32.7 83.0 27.3 30.0 72.0 62.0 28.7 23.0
Qwen2.5-VL-32B [2] 45.6 87.2 40.7 83.0 29.5 40.7 74.0 60.0 27.3 29.5
Qwen2.5-VL-72B [2] 43.7 90.8 34.0 87.0 27.9 30.6 74.0 61.0 26.7 25.5
LLaVA-OneVision-7B [23] 43.7 80.0 40.0 56.0 31.7 39.4 65.0 33.0 38.0 32.0
LLaVA-OneVision-72B [23] 48.7 79.2 50.7 67.0 37.9 40.0 76.0 41.0 38.7 39.3
InternVL3-8B [77] 47.9 83.6 49.3 75.0 34.5 36.9 70.0 44.0 40.0 37.0
InternVL3-38B [77] 51.0 85.6 56.0 71.0 42.6 40.0 77.0 45.0 47.3 35.0
InternVL3-78B [77] 52.3 87.6 54.7 77.0 42.6 36.6 76.0 56.0 46.0 40.3
Visual Grounded Reasoning Models
Pixel-Reasoner-7B [50] 49.7 89.6 52.0 86.0 38.9 30.9 71.0 72.0 46.0 32.5
DeepEyes-7B [76] 53.2 90.0 52.7 89.0 43.3 33.4 76.0 69.0 44.0 35.0
TreeVGR-7B 54.9 87.6 50.7 83.0 47.0 43.4 74.0 66.0 51.3 39.0
Table 4: Comparison with state-of-the-art alternatives on MME-RealWorld-Lite [72]. All results are self-collected. The best performance is highlighted in bold.
@@
Correlation with Other Multimodal Benchmarks. We systematically compare our TreeBench with three existing multimodal benchmarks: AI2D [17], MathVista [37], and MMStar [7], in Figure 13, Figure 13, and Figure 13, respectively, to investigate potential performance correlations. Our analysis reveals a decoupling of performance characteristics. For instance, while GPT-4o-1120 [41] ranks among the top performers on TreeBench, it lags significantly behind alternatives on other benchmarks. This dissociation underscores the unique emphasis on “thinking with images” of our TreeBench.
Capability Benchmark Qwen2.5-VL-7B [2] TreeVGR-7B Qwen2.5-VL-72B [2]
Vision-centric question answering CV-Bench-2D [53] 74.1 76.9 $\uparrow$ 2.8 77.7
CV-Bench-3D [53] 72.6 77.6 $\uparrow$ 5.0 87.0
MMVP [54] 66.7 75.3 $\uparrow$ 8.6 66.7
@@
POPE [26] 86.7 87.2 $\uparrow$ 0.5 84.9
HallusionBench [14] 48.2 50.1 $\uparrow$ 1.9 55.2
@@
Table 5: Comparison with state-of-the-art alternatives on other multimodal benchmarks. Results are obtained from [16], otherwise are self-collected.
5.2 Main Properties of TreeVGR
Baselines and Benchmarks. We mainly compare our TreeVGR with state-of-the-art private models, including GPT-4o-1120 [41] and o3-0416 [43] from OpenAI, and Gemini-2.5-Flash-0520 [9] and Gemini-2.5-Pro-0605 [10] from Google DeepMind. Additionally, 8 representative open-source general models are incorporated, including LLaVA-OneVision series [23], Qwen2.5-VL series [2], and InternVL3 series [77]. Furthermore, two very recent visual grounded reasoning models are also included, i.e., DeepEyes [76] and Pixel-Reasoner [50], as both of them follow a “grounding then answering” pipeline, with the capability of “thinking with images”. Evaluations are mainly conducted on TreeBench, V* Bench [66], HR-Bench [63], and MME-RealWorld-Lite [72].
Results on High-Resolution Benchmarks. As demonstrated in Table 2, our TreeVGR-7B achieves 45.2 on our TreeBench, even surpasses Qwen2.5-VL-72B [2], which is comparable with open-source state-of-the-art InternVL3-78B [77]. Moreover, compared with visual grounded reasoning models, our TreeVGR not only achieves a higher overall performance, but also obtains larger mIoU, indicating its effectiveness in precisely localizing target objects. Additionally, in Table 3, TreeVGR also achieves open-source state-of-the-art on V* Bench [66] with 94.3 accuracy, which is even comparable with o3 [43]. On HR-Bench [63] and MME-RealWorld-Lite [72] illustrated in Table 3 and Table 4, respectively, our TreeVGR brings significant improvements over our base model, Qwen2.5-VL-7B [2].
@@ 5.2 Main Properties of TreeVGR
Results on Other Multimodal Benchmarks. In Table 5, we compare our TreeVGR with its base model Qwen2.5-VL-7B [2] on a variety of conventional multimodal benchmarks. Specifically, we select CV-Bench [53], MMVP [54], and RealWorldQA [68], to evaluate vision-centric question-answering capabilities. MMBench [30], POPE [26], and HallusionBench [14] are selected for evaluating general VQA capabilities, and AI2D [17] and ChartQA [38] for comprehension with document and chart. We observe significant improvements in most cases, especially for vision-centric benchmarks. Notably, TreeVGR-7B achieves 75.3 on MMVP [54], even surpasses Qwen2.5-VL-72B [2] by a significant margin.
Rewards TreeBench V* [66] MME-RW [72]
@@
Table 6: Ablations of each component of our TreeVGR. “MME-RW” stands for MME-RealWorld-Lite [72], and “Acc” represents the multiple-choice accuracy. This improvement mainly comes from the training set, as many training samples from V* [66] are included in RL. The model tends to enumerate candidate boxes to obtain larger IoU recall, and fails to produce final answers.
5.3 Ablation Studies
@@
6 Conclusion
TreeBench addresses critical gaps in existing benchmarks by focusing on three principles: focused visual perception (identifying subtle targets in cluttered scenes), traceable evidence (quantifiable reasoning chains via bounding box annotations), and vision-centric second-order reasoning. Constructed through expert-driven annotation and multi-stage quality control, TreeBench features 405 high-difficulty visual question-answer pairs with precise bounding boxes, emphasizing small objects in real-world scenarios. It reveals the limitations of state-of-the-art models, e.g., OpenAI-o3 [43] scores 54.8%, while setting a new standard for assessing nuanced visual grounding, multi-step reasoning transparency, and cross-modal interaction.
TreeVGR advances VGR training through reinforcement learning guided by dual IoU rewards, which explicitly supervise bounding box generation to ensure both precision and recall. This approach enables explainable reasoning pathways and achieves significant improvements across benchmarks. TreeVGR achieves superior performance in just 5 epochs compared to 32 epochs for [76].

### v2

@@ Appendix C Benchmark Statistics
Figure 11: Performance decoupling with AI2D (Hiippala et al., 2021).
Figure 12: Performance decoupling with MathVista (Lu et al., 2023).
Figure 13: Performance decoupling with MMStar (Chen et al., 2024a).
Appendix D Analysis of TreeBench
Correlation between Localization and Performance. Importantly, for visual grounded reasoning models, our traceable evaluation demonstrates a positive correlation between localization precision and the overall performance, as illustrated in Table 2. This positive correlation between precise localization (mIoU) and overall performance is evident in the progressive improvement from DeepEyes-7B (Zheng et al., 2025b) to Pixel-Reasoner-7B (Su et al., 2025) to our final TreeVGR-7B. As mIoU increases, the overall scores rise correspondingly, with TreeVGR-7B achieving the highest mIoU and strongest overall performance at the same time.
@@ Appendix D Analysis of TreeBench
Correlation with Other Multimodal Benchmarks. We systematically compare our TreeBench with three existing multimodal benchmarks: AI2D (Hiippala et al., 2021), MathVista (Lu et al., 2023), and MMStar (Chen et al., 2024a), in Figure 13, Figure 13, and Figure 13, respectively, to investigate potential performance correlations. Our analysis reveals a decoupling of performance characteristics. For instance, while GPT-4o-1120 (OpenAI, 2024a) ranks among the top performers on TreeBench, it lags significantly behind alternatives on other benchmarks. This dissociation underscores the unique emphasis on “thinking with images” of our TreeBench.
The Quality of Visual Evidence in TreeBench. First, in Table 5, we mask all instances during inference on TreeBench. The results show a significant performance drop across all models. This confirms that the bounding boxes in TreeBench are not only high-quality but also indispensable for accurate visual grounded reasoning, directly validating the importance of the annotated evidence. Moreover, we conduct dual experiments in Table 6. When we provide ground-truth bounding boxes as explicit evidence hints to models, all models achieve consistent performance gains. It indicates that bounding boxes indeed help models explicitly anchor their reasoning to visual evidence: without boxes, models may rely on ambiguous textual biases or global image impressions, while with boxes, they are forced to align their answers with the specific visual content in the evidence region, which reflects a shift from “heuristic reasoning” to “evidence-based reasoning”.
Table 5: Performance comparison with masked target instances. When masking out all target instances on TreeBench, we observe a significant performance drop across all models, confirming that the annotated bounding boxes are not only high-quality but also indispensable for accurate visual grounded reasoning.
Masking Qwen2.5-VL-7B InternVL3-8B GPT-4o o3 Gemini-2.5-Flash Gemini-2.5-Pro
37.0 38.8 46.9 54.8 45.9 54.1
31.8 $\downarrow$ 5.2 29.6 $\downarrow$ 9.2 29.1 $\downarrow$ 17.8 33.8 $\downarrow$ 21.0 29.9 $\downarrow$ 16.0 33.1 $\downarrow$ 21.0
Table 6: Performance comparison with explicit bounding boxes-based textual hints. When we provide ground-truth bounding boxes as explicit evidence hints to models, all models achieve consistent performance gains.
Textual Boxes Qwen2.5-VL-7B InternVL3-8B GPT-4o o3 Gemini-2.5-Flash Gemini-2.5-Pro
37.0 38.8 46.9 54.8 45.9 54.1
43.7 $\uparrow$ 6.7 43.5 $\uparrow$ 4.7 49.4 $\uparrow$ 2.5 58.3 $\uparrow$ 3.5 51.9 $\uparrow$ 6.0 61.0 $\uparrow$ 6.9
Table 7: Comparison with state-of-the-art alternatives on other multimodal benchmarks, including CV-Bench (Tong et al., 2024a), MMVP (Tong et al., 2024b), MMBench (Liu et al., 2023b), POPE (Li et al., 2023c), AI2D (Hiippala et al., 2021), and ChartQA (Masry et al., 2022). Results are obtained from (Guo et al., 2025b), otherwise are self-collected.
Capability Benchmark Qwen2.5-VL-7B TreeVGR-7B Qwen2.5-VL-72B
Vision-centric question answering CV-Bench-2D 74.1 76.9 $\uparrow$ 2.8 77.7
CV-Bench-3D 72.6 77.6 $\uparrow$ 5.0 87.0
MMVP 66.7 75.3 $\uparrow$ 8.6 66.7
@@
POPE 86.7 87.2 $\uparrow$ 0.5 84.9
@@
Appendix E Implementation Details
E.1 Cold-Start Initialization
Data Construction. We base our supervised fine-tuning (SFT) dataset on VGR-158K (Wang et al., 2025e), which provides pseudo-chain-of-thought annotations paired with bounding boxes for visual reasoning tasks. However, to align with the grounding capabilities of our base model (Qwen2.5-VL series (Bai et al., 2025a)), which outputs absolute coordinates rather than the normalized coordinates (ranging from 0 to 1) used by LLaVA-NeXT (Liu et al., 2024a) in (Wang et al., 2025e), we perform coordinate system conversion. Specifically, for each bounding box, we transform normalized coordinates $r_{x_{1}},r_{y_{1}},r_{x_{2}},r_{y_{2}}]$ into absolute coordinates via $x_{1},y_{1},x_{2},y_{2}]=[Wr_{x_{1}},Hr_{y_{1}},Wr_{x_{2}},Hr_{y_{2}}]$, where $H\times W$ is the resolution of the input image. Next, we filter samples to prioritize complex reasoning pathways, retaining only entries with multiple bounding boxes (i.e., more than one box per reasoning trajectory). This yields 35K samples, as multi-box interactions demand stronger spatial-temporal reasoning compared to single-box tasks. Subsequently, we construct a reflective subset of 4.7K samples among them by introducing controlled perturbations: for each sample, we (1) inject a synthetic error by inserting a randomly generated incorrect bounding box into the reasoning sequence, and (2) append the meta-cognitive prompt “Wait, this box seems to be wrong” immediately afterward, resulting in our TreeVGR-SFT-35K. This design explicitly trains the model to detect and correct erroneous visual grounding, which is a critical skill for robust real-world deployment.
Optimization. Initialized from Qwen2.5-VL-7B-Instruct (Bai et al., 2025a), we train TreeVGR-7B-CI (“CI” here stands for Cold Initialization) with 8 GPUs using LLaMA-Factory (Zheng et al., 2024), where the AdamW optimizer (Loshchilov and Hutter, 2017) with a learning rate of 5e-6 and a global batch size of 256 is utilized. The learning rate is decayed following a cosine schedule (Loshchilov and Hutter, 2016) with a warmup ratio of 0.1.
E.2 Reinforcement Learning
Data Construction. TreeVGR incorporates a novel dual IoU reward, which means each sample should contain ground-truth bounding boxes during the RL phase. To this end, we filter hard samples from the original 191K training set of V* (Wu and Xie, 2024) using Qwen2.5-VL-7B-Instruct (Bai et al., 2025a), resulting in 30K samples. Additionally, we incorporate the VisDrone dataset (Zhu et al., 2021), which is originally designed for detection and tracking under UAV images, which offers extremely high-resolution, real-world scenes with a large number of small and varied objects and their corresponding bounding box annotations. We reformulate the training set and the validation set into 38K multiple-choice counting problems, and only retain samples with the ground-truth number ranging from 5 to 10, contributing to the final 7K samples. Finally, our TreeVGR-RL-37K consists of 30K open-ended question-answering samples from V* (Wu and Xie, 2024) and 7K multiple-choice problems from VisDrone (Zhu et al., 2021).
Optimization. Initialized from TreeVGR-7B-CI, we train our final TreeVGR-7B with 8 GPUs, with another 8 GPUs serving the reward model, i.e., Qwen2.5-72B-Instruct (Team, 2024), using vLLM (Kwon et al., 2023). We adopt Group Relative Policy Optimization (GRPO) (Shao et al., 2024b), which has been proved to be effective and efficient for diverse tasks. We have also tried DAPO (Yu et al., 2025), but we find it unstable compared with GRPO. Therefore, we simply utilize the original GRPO (Shao et al., 2024b). We implement using EasyR1 (Zheng et al., 2025a), which is a clean fork of veRL (Sheng et al., 2024). We train our TreeVGR-7B with 5 epochs on TreeVGR-RL-37K, which is significantly less than DeepEyes-7B (Zheng et al., 2025b) (which is trained on 47K samples with 32 epochs).
Appendix F More Experiments
F.1 Results on Other Multimodal Benchmarks
@@
In Table 7, we compare our TreeVGR with its base model Qwen2.5-VL-7B (Bai et al., 2025a) on a variety of conventional multimodal benchmarks. Specifically, we select CV-Bench (Tong et al., 2024a) and MMVP (Tong et al., 2024b) to evaluate vision-centric question-answering capabilities. MMBench (Liu et al., 2023b) and POPE (Li et al., 2023c) are selected for evaluating general VQA capabilities, and AI2D (Hiippala et al., 2021) and ChartQA (Masry et al., 2022) for comprehension with document and chart. We observe significant improvements in most cases, especially for vision-centric benchmarks. Notably, TreeVGR-7B achieves 75.3 on MMVP (Tong et al., 2024b), even surpasses Qwen2.5-VL-72B (Bai et al., 2025a) by a significant margin.
Table 8: Ablations of each component of our TreeVGR. “MME-RW” stands for MME-RealWorld-Lite (Zhang et al., 2024a), and “Acc” represents the multiple-choice accuracy. This improvement mainly comes from the training set, as many training samples from V* (Wu and Xie, 2024) are included in RL. The model enumerates boxes to obtain larger IoU recall, and fails to produce final answers.
Rewards TreeBench V* MME-RW
@@
F.2 Ablation Studies
@@
[…]
Appendix G Limitations and Future Works
One possible limitation of TreeVGR is the model scale and architecture, which is limited to Qwen2.5-VL-7B (Bai et al., 2025a). Experiments with other base models and larger model scales could be future work. Furthermore, TreeVGR is not a general multimodal reasoner, as it is not designed to perform ultra-long reasoning processes in math, sciences, and coding. How to effectively unify vision-centric reasoning models with standard text-centric models could be a future work.
As for TreeBench, we find that the “perspective transform” protocol becomes one of the major bottlenecks, which means current state-of-the-art multimodal models, even including visual grounded reasoning models, have not effectively modeled the ego-view 3D awareness. Moreover, basic perception capabilities under complex scenes are also limited, leading to relatively low scores on “attributes” and “material”. How to effectively let LMMs perceive any details of the dense visual world becomes a critical challenge.
Appendix H Qualitative Examples of TreeVGR
We first provide qualitative examples of the correlation between the predicted bounding boxes and the model’s internal attention maps in Figure 15, where the attention values are the mean attention of generated tokens related to bounding boxes (highlighted in red). Strong correlations are observed across examples, indicating that the model’s focus is truly guided by those textual bounding boxes-based hints. This explains why visual grounded reasoning pipelines could be such effective. Note that sometimes the visual attention is still being overwhelmed by textual domain knowledge even if the grounding is accurate (see the failure case in Figure 15). Explicitly enforcing the model’s focus beyond attention may be one potential solution.
Next, we provide qualitative examples and failure cases of our TreeVGR on each category of TreeBench in Figures 16 to 25.
Figure 15: Qualitative examples (first three rows) and failure cases (the last row) on V* Bench (Wu and Xie, 2024) with attention maps.
Figure 16: Qualitative examples (left) and failure cases (right) on the “Attributes” protocol of TreeBench.
Figure 17: Qualitative examples (left) and failure cases (right) on the “Material” protocol of TreeBench.
Figure 18: Qualitative examples on the “Physical State” protocol of TreeBench.
Figure 19: Qualitative examples (left) and failure cases (right) on the “Object Retrieval” protocol of TreeBench.
Figure 20: Qualitative examples on the “OCR-Integrated Question-Answering” protocol of TreeBench.
Figure 21: Qualitative examples (left) and failure cases (right) on the “Perspective Transform” protocol of TreeBench.
Figure 22: Qualitative examples (left) and failure cases (right) on the “Ordering” protocol of TreeBench.
Figure 23: Qualitative examples (left) and failure cases (right) on the “Contact and Occlusion” protocol of TreeBench.
Figure 24: Qualitative examples (left) and failure cases (right) on the “Spatial Containment” protocol of TreeBench.
Figure 25: Qualitative examples (left) and failure cases (right) on the “Comparison” protocol of TreeBench.